OpenStudy (btsangiie13280):

A cheetah can travel up to 29 meters per second. What is the cheetah’s speed, to the nearest tenth, in miles per hour? (1 mile ≈ 1609 meters)

OpenStudy (luckycoins888):

wow

OpenStudy (luckycoins888):

In the middle of solving it my brain failed

OpenStudy (luckycoins888):

-_-

OpenStudy (btsangiie13280):

lol i have this due in 3 hours i have more problems to go, i swear my teachers are killing me :'(

OpenStudy (jclark):

It's a proportion. a/b = c/d

OpenStudy (jclark):

Meters/sec = miles /hour. Which is the unknown. Call it x

OpenStudy (mrnood):

there are 3600 seconds in 1 hour so if it can run 29m in 1 second how many METRES can it run in 1 hour? That will give you a speed in "Metres per Hour" You then divide THAT by 1609 to give you miles per hour

OpenStudy (jiteshmeghwal9):

\[29 m/s = x miles/hr\] 1m= 1/1609 miles 1s=1/3600 hour

OpenStudy (jiteshmeghwal9):

\[\frac{29 \times \frac{1}{1609}}{\frac{1}{3600}} miles / hour = x \space miles/ hour\]